STATEMENT OF PURPOSE

I would like to earn Ph.D. degree in Computer Engineering. I want to improve my knowledge about the subjects in Computer Science and Engineering since my purpose is to become a faculty member and also a researcher in this field.

I am interested in Artifical Intelligence (AI) methods (Expert Systems, Artificial Neural Networks and Genetic Algorithms). And also, I want to take some technical and practical courses especially Mathematical Modelling, Optimization Theory, Statistical Models, Signal Processing, System Theory, Information Theory, Computer Vision, Pattern Recognition, Modelling and Performance Evaluation of Computer System.

I was born on January 18th, 1969 in Erzurum, Turkey. I graduated from Hacettepe University, Deparment of Computer Science and Engineering in 1991. I got my M.Sc. degree related to Image Processing in BioMedical Applications in 1995. I worked as a research and teaching assistant between 1992 and 1995 at the same department. Through the second year of undergraduate to research assistantship, I had an oppurtinity to have experience on some operating systems such as DOS, AOS/VS, UNIX and to use some pogramming languages such as Pascal, C, C++, Smalltalk, PL/I, COBOL, Fortran77, BASIC, INTEL 8080A/8085, INTEL 8086/8088, etc. I have very good experience about database systems such as SQL, ORACLE, and Progress.

As an undergraduate project, I worked with Prof. M. Umit Karakas for the MYDEARBABY project. This is an expert system that monitorizes Fetal Heart Rate.

In the first year of my graduate study, I prepared a technical report about Genetic Algorithms and applications.

As a final exam of the graduate Artificial Neural Network(ANN) course, I implemented a risk evalution system for loaning money to customers. After traning, this system evaluates each customer’s risk for loaning him/her money by using his/her personel profile (with information on his/her demanded loan). I applied the Back Propagation learning mechanism in the this risk evaluation system.

In the graduate Microprocessor course, I implemented a Local Area Network card and a software that controls the communication between four PCs, and between PCs and one printer by using this card. This card is placed into extention slot of a dedicated personal computer (IBM/AT). The other PCs are connected via RS232 cable to the card and the printer is connected to the card with a parallel interface component. The software for this system was written in C programming language.

As a M.Sc. thesis, I implemented an Image Processing system for Fetal Head Images. The aim of this study is to develop and to implement a system to process fetal ultrasound images. The sytem has two main parts : the first part is a menu driven image processing tool for ultrasound images, and the second part is an automated measurement on these fetal head images. The first part of the system includes many image processing techniques such as enhancement, segmentation, and texture analysis. In the second part, the ultrasonographic parameters related to fetal head such as the Biparietal diameter (BPD) and the Head circumference (HC) are evaluated automatically from the image.

I have publications about my graduate studies, these are :

Ozdemir, H.T., Odcikin, Z., Karakas, U., Beksac, M.S., Cakar, N.; A Digital Image Processing Subsytem for Semi Automated Karyotyping; International Symp. on Computer and Information Science, November 7-9, 1994, Antalya, Turkey.

Odcikin, Z., Ozdemir, H.T., Karakas, U., Beksac, M.S.; Design and Implementation of Image Processing Sytem for Fetal Head Images; 11th Informatique Conf. of Turkey, Sept. 14-17, 1994, Istanbul, Turkey (in Turkish).

I worked as a research/teaching assistant at Hacettepe University, Department of Computer Science and Engineering for the last three years. During this period, I was responsible for Database Laboratory courses, Digital Systems Laboratory courses, Microprocessing Laboratory courses, and Software Laboratory courses for undergraduate students.

I have been living in Syracuse since January, 1995 because of my husband’s graduate study at Syracuse University. I have no problem with living expenses. Therefore, I am only looking for a financial support for the tuition and fee expenses. Since I have a three year experience in assistantship and I believe that I am qualified for graduate assistantship, I am applying to the graduate assistantship or scholarship to pay the tuition and fee.

Zeynep Odcikin OZDEMIR